OpenStudy (anonymous):

WILL FAN AND MEDAL!!! Which circles have an area of 9pi m2? Choose all answers that are correct. A. a circle with a radius of 3 m B. a circle with a diameter of 6 m C. a circle with a radius of 9 m D. a circle with a diameter of 9 m

Directrix (directrix):

The area of a circle is pi * r^2. To help you get started checking options, A) a circle with a radius of 3 m Area = pi * 3^2 = 9 pi square meters. Therefore, A is one of the correct answers.

Directrix (directrix):

@abelwell Now, you check this one, okay? B. a circle with a diameter of 6 m

OpenStudy (anonymous):

i still dont get it. (i SUCK at math) could you explain it a bit more? @Directrix

Directrix (directrix):

On option B, the diameter is 6 so the radius is 3. In the formula, replace r by 3 and see what you get for the area. A = pi * r^2

OpenStudy (anonymous):

9? (sorry if its wrong :/)

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

9 is correct if you add the pi, so 9 pi. A = pi * r^2 ------------ That means that Option B is correct.

Directrix (directrix):

On option C, the radius is 9 and when squared, it will be 81 along with the pi which is not correct.

Directrix (directrix):

On option D, the radius is 9/2 or 4.5 which when squared will not be 9.

Directrix (directrix):

@abelwell The two correct answers are what? Post what you think, okay?

OpenStudy (anonymous):

A and B?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Correct.